### v1

# Denis Roio Dyne.org &Rebecca Selvaggini

Abstract
It is of critical importance to design digital identity systems that ensure the privacy of citizens as well protect them from issuer corruption. Unfortunately, what Europe’s and USA’s public sectors are currently developing does not offer such basic protections. Aiming at solving this issue, we propose a method for untraceable selective disclosure and privacy preserving revocation of digital credentials, utilizing the unique homomorphic characteristics of second order Elliptic Curves and Boneh-Lynn-Shacham (BLS) signatures operated on them. Our approach ensures that users can selectively reveal only the necessary credentials, while protecting their privacy across multiple presentations. We also aim to protect users from issuer corruption, by making it possible to apply a threshold for revocation to require collective agreement among multiple revocation issuers.
1 Introduction
Digital identity systems implement credential issuance and presentation mechanisms so that a person (holder) can voluntarily disclose his or her own acquired skills, professed attributes, or completed accomplishments. Credentials are signed by issuer authorities and encapsulated within various forms of digital proofs to be held in digital wallets, empowering individuals to reveal only chosen details to designated recipients, to limit data exposure and permit a user-controlled release of information.
Such systems are known as selective disclosure and this article aims at improving their cryptography to adhere to basic privacy-by-design standards.
1.1 State of the art
Selective disclosures are being used by nation states across the world in their next generation identity systems, for instance EIDAS2.0 in Europe where the European Digital Identity Wallet Architecture and Reference Framework[1] mandates the use of SD-JWT[2] and mDOC [3]. Unfortunately both SD-JWT and mDOC adopt for their cryptography just Hash-Based Message Authentication Codes (HMAC) to generate proofs, and they can be easily traced (non-unlinkability): recent critical feedback on the EUDI ARF[4] details several problems.
In North America the efforts concentrate on the adoption of the the BBS+ algorithm[5] leveraging its Zero Knowledge Proof properties and applied to W3C Verifiable Credentials[6] to obtain an higher degree of privacy.
1.2 Threats considered
The different choices in data formats in these two approaches is irrelevant in relation with cryptography, any choice between JSON Web Tokens or W3C Verifiable Credentials does not impact the privacy level. But the cryptography adopted determines the adequacy of a solution to face three important threats that can render an algorithm unsuitable to be used in real world situations.
Linkability
The EUDI-ARF standard dictates that credentials issued to a holder, can be presented (in the form of a verifiable presentation) to a relying party in order to have one or more attributes verified. Every verifiable presentation includes one or more HMAC(s), formatted in SD-JWT: the HMACs are identical each time a verifiable presentation is produced from a certain credential. This makes possible for colluding relying parties, or to malicious actors, to trace a holder’s identity by collecting, exchanging and confronting verifiable presentations. In order to preserve the privacy of a credential holder the disclosures presented by the wallet should not be traceable across different presentations (unlinkability). This threat appears to be well mitigated by BBS+ through its Zero Knowledge Proof implementation.
Privacy breach of revocation lists
There is no privacy-preserving revocation system designed, either in EUDI-ARF, or in W3C-VC and BBS+. The approach followed for the revocation is that of the status (revocation) lists [7] for W3C-VC and EUDI-ARF is investigating a similar method. The use of a certificate status lists (CRL) presents issues related primarily to privacy [8], because sensitive information about holders leaks from the list. This is partially mitigated by expiration dates, in cases where credentials can be short-lived (typically less important credentials), but not applicable with digital identification documents such as ID, driving license, passport and social security numbers, which typically have longer or no expiration time. In case the choice of strategy for revocation is left open to developers, the risk for major privacy breaches may occur, for example with the adoption of public status lists [9]. Based on these arguments, we believe that anonymous revocations are a sine qua none condition to guarantee a sufficient level of privacy in digital identities and credentials. Our approach is that of designing a privacy-preserving revocation mechanism to remove the leak of holder’s information.
Revocation issuer corruption
If the choice of interactive revocation is left to a single issuer, one may unilaterally choose to revoke credentials, without being subject to revision or having to seek consensus with a quorum of issuers. This situation leads to security issues in case Issuers are corrupted and make a weaponized use of digital revocations to persecute engaged individuals. Such a condition becomes a real concern for journalists or activists living under dictatorial regimes that may arbitrarily revoke their credentials, or even ID cards and passports. Similarly, a security breach of an issuer service, would result in similar threats. We mitigate this risk by introducing the possibility for threshold issuance of revocation keys.
2 Overview
2.1 Key contributions
The cryptographic scheme described in the paper, name SD-BLS for brevity, takes inspiration by modern zero knowledge proofs schemes such as [5]BBS signatures or Coconut[10]. The paper aims to employ some of the properties of the said schemes, overcome some of the shortcomings of SD-JWT and mDOC formats, mainly the non-unlinkability of credentials, proposing a novel cryptographic approach to similar data structures. Furthermore SD-BLS proposes novel anonymous cryptographic revocation flow for verifiable credentials, that aims to solve the privacy issues posed by status and revocation lists.
Unlinkabe selective disclosure
Similarly to the SD-JWT and mDOC formats, SD-BLS produces an array of claims: the elements of the array are individually signed by the issuer. In SD-BLS the signature(s) replace the HMAC. Using the homomorphic properties of this curve SD-BLS enables the holder to produce a different presentation each time they want to prove a claim. SD-BLS can be used to prove credentials in public (on blockchain) when they are boolean attributes.
Anonymous cryptographic revocation
SD-BLS proposes a novel approach to credential revocation: the data published by the revocation issuer will produce cryptographic material that contains no information about and is never linkable to the holder. The cryptographic revocation material allows anyone to verify if an SD-BLS proof produced by a credential holder has been revoked by one or multiple issuers. The unlinkability, and thus anonymity of the cryptographic revocation, allows the revocation issuer to publish revocations off or on blockchain.
2.2 Applications
In this section we present some applications and use cases for digital identity and credentials, that could benefit from using the unlinkability and anonymous revocation of the the SD-BLS scheme.
Digital identity
The focus of the EUDI-ARF specifications is identity documents: it defines mechanisms and data structures to issue a Personal Identification (PID) as well as digital driving licenses. Similarly, the US government is experimenting with W3C-VC and mDOC for cross-states interoperable driving licenses. SD-BLS data format is similar to SD-JWT and mDOC, offering selective disclosure and adding unlinkability and anonymous revocation.
Academic credentials
Diploma and academic credentials are among the core offerings of EBSI as well as a primary research target of the W3C VC working group.
KYC/AML
We are unaware of standardization efforts for interoperable credentials in the fields of "Know Your Customer" (KYC) and Anti Money-Laundering (AML) certifications. We are aware of solution providers experimenting with W3C-VC for AML applications.
Generic light credentials
As the digital identity and verifiable credential technologies are maturing, they are being considered for usage in less privacy concerning applications, such subscriptions and membership and fidelity cards.
Verifiable credentials on Blockchain
SD-BLS can be used on blockchain to produce publicly verified proofs when the credentials are made of clear-text messages stating a boolean property or ACL clearance.
  • A verifiable presentation can be used by a smart-contract to verify if one or more holder’s claim match the requirement needed to process a transaction. The verifiable presentation is anonymous, therefore its storage in the transaction log doesn’t pose privacy issues and is GDPR compatible.
  • Issuers can publish their cryptographic revocation lists on chain, allowing smart-contracts to verify the status of a credential, still granting unlinkability.
  • The issuer’s public keys can also be published on-chain, although this does not represent a novelty.
It’s worth mentioning that, while the verifiable presentation is itself unlinkable to the holder’s blockchain address, the unlinkability of the holder to their blockchain address at verification phase, needs to be addressed within the blockchain protocol used: this issue is out of scope in this paper.
3 Implementation
In this section we will provide a detailed description of the algorithm we propose for selective disclosure and unlinkable revocation using BLS signatures.
3.1 Notations and assumptions
We will adopt the following notations:
  • $\mathbb{F}_{p}$ is the prime finite field with $p$ elements (i.e. of prime order $p$);
  • $E$ denotes the (additive) group of points of the curve BLS12-381 [11] which can be described with the Weierstrass form $y^{2}=x^{3}+16$;
  • $E_{T}$ represents instead the group of points of the twisted curve of BLS12-381, with embedding degree $k=12$. The order of this group is the same of that of $E$;
We also require defining the notion of a cryptographic pairing. That is defined as a function $e:\mathbb{G}_{2}\times\mathbb{G}_{1}\to\mathbb{G}_{T}$, where $\mathbb{G}_{1},\mathbb{G}_{2}$ and $\mathbb{G}_{T}$ are all groups of same order $n$, such that satisfies the following properties:
  • i.
    Bilinearity, i.e. given $P_{1},Q_{1}\in\mathbb{G}_{1}$ and $P_{2},Q_{2}\in\mathbb{G}_{2}$, we have
    $e(P_{2},P_{1}+Q_{1})=e(P_{2},P_{1})\cdot e(P_{2},Q_{1})$
    $e(P_{2}+Q_{2},P_{1})=e(P_{2},P_{1})\cdot e(Q_{2},P_{1})$
  • ii.
    Non-degeneracy, meaning that for all $g_{1}\in\mathbb{G}_{1},g_{2}\in\mathbb{G}_{2}$, $e(g_{2},g_{1})\neq 1_{\mathbb{G}_{T}}$, the identity element of the group $\mathbb{G}_{T}$;
  • iii.
    Efficiency, so that the map $e$ is easy to compute;
  • iv.
    $\mathbb{G}_{1}\neq\mathbb{G}_{2}$, and moreover, that there exist no efficient homomorphism between $\mathbb{G}_{1}$ and $\mathbb{G}_{2}$.
For the purpose of our protocol we will have $\mathbb{G}_{1}=E$ and $\mathbb{G}_{2}=E_{T}$, and $\mathbb{G}_{T}\subset\mathbb{F}_{p^{12}}$ is the subgroup containing the $n$-th roots of unity, where $n$ is the order of the groups $E$ and $E_{T}$. Instead $e:E_{T}\times E\to\mathbb{G}_{T}$ is the Miller pairing, which in our work is encoded as the method miller(ECP2 P, ECP Q).
3.2 Issuance
As for other well known algorithms BLS signing will work following three main steps:
  • Key Generation phase. For an issuer who wants to sign a credential $m$, a secret key $sk$ is a random number chosen uniformly in $\mathbb{F}_{n}$, where $n$ is the order of the groups $\mathbb{G}_{1},\mathbb{G}_{2},\mathbb{G}_{T}$. The corresponding public key $pk$ is the element $sk\cdot G_{2}\in E_{T}$;
  • Signing phase. The credential $m$ is first hashed into the point $U\in E$, which in our scheme is done by the method hashtopoint; the related signature is then given by $\sigma=sk\cdot U$;
  • Verification phase. For an other user that wants to verify the authenticity and the integrity of the message $m$, it needs to
    • 1.
      parse $m,pk$ and $\sigma$
    • 2.
      hash the message $m$ into the point $U$ and then check if the following identity holds,
      $e(pk,U)=e(G_{2},\sigma)$
    If verification passes it means that $\sigma$ is a valid signature for $m$.
  • Proof of correctness By using the definitions of the elements involved and exploiting the property of the pairing $e$ we have
    $\begin{split}e(pk,U)&=e(sk\cdot G_{2},U)\\ &=e(G_{2},U)^{sk}\\ &=e(G_{2},sk\cdot U)\\ &=e(G_{2},\sigma)\end{split}$
BLS signatures also support aggregation: it is possible to aggregate a collection of multiple signatures $\sigma_{i}$ (each one related to a different message $m_{i}$) into a singular new object $\sigma$, that can be validated using the respective public keys $pk_{i}$ in a suitable way.
Since $\sigma_{i}\in G_{1}\forall i$, the algorithm has an homomorphic property. We exploit this property to add a revocation signature into the signed credential.
We create a new secret key $rev$ with public key $r_{2}$, then compute the point $r_{1}\in\mathbb{G}_{1}$ and signature of the claim $\sigma_{rev}$ to add to the issuer signature:
$\begin{split}r_{2}&=rev\cdot G_{2}\\ r_{1}&=rev\cdot G_{1}\\ \sigma_{rev}&=rev\cdot U\\ \sigma&=sk\cdot U+\sigma_{rev}\\ \end{split}$
The set of all the signed claims will be:
$\mathcal{C}=\big{\{}\{r_{1i},r_{2i},\sigma_{i}\}:i\in claims\big{\}}$
At the end of this phase the holder is sent the signed claims to be stored in a private wallet, while the issuer stores a list of revocations, associated with an identifier of the holder and the specific credential, into a private database that can be used later to issue revocations.
3.3 Presentation
As we have done in the issuance phase (3.2), we exploit the homomorphic property of BLS signatures once again, but this time to render the presented signature unlinkable (impossible to trace across multiple presentations) by adding a "blinding factor" signature to the one provided by the issuer.
Let $\mathcal{H}$ denote a cryptographic hash function (e.g. SHA-256) and let $\varepsilon$ be an integer chosen uniformly at random that serves as the blinding factor. Using the same technique of the Tripartite Diffie-Hellman Key Exchange [12] we construct a secret that is binded to the blinding factor, the Issuer public key and the "revocation point" $r_{1}=rev\cdot G_{1}$ computed in the issuance phase:
$\tau=\mathcal{H}(e(A.pk,r_{1})^{\varepsilon})$ (1)
When the holder needs to disclose a particular claim $m$ we construct the credential proof as follow: let $\sigma$ be the signature of the claim provided by the Issuer, then apply the blinding factor to the signature using the secret $\tau$
$s=\sigma+sign(\tau,m)$ (2)
and the corresponding public key computed as
$p=r_{2}+G_{2}\cdot\tau$
where $r_{2}$ is the revocation key linked to the claim. We also attach the public key $r=G_{1}\cdot\varepsilon$ that correspond to the secret key $\varepsilon$.
3.4 Verification
Credential verification is made by checking that the issuer has signed the blinded proof being presented. In order to do that the credential issuer’s public key must be added to the proof.
As explained in section 3.3 we can consider the credential proof as a collection of tables of the following form:
$c_{p}=\{id,p,r,s\}$
where $s$ and $p$ are respectively the signature and the public key of the credential proof of the string $id$ after verifying the status of the revocation, we can check the validity of the presented claim computing the key
$pk=A.pk+p$
and verify the bls signature.
3.4.1 Proof of correctness
The signature $s$ is given by
$\begin{split}s&=\sigma+sign(\tau,id)\quad\text{where}\\ \sigma&=sign(A.sk,id)+sign(rev,id)\\ \text{so}\quad s&=A.skU+revU+\tau U\end{split}$
where $U$ is the mapping of the string $id$ in the group $G_{1}$.
Recalling the verification formula, it holds that:
$\begin{split}e(pk,U)&=e(A.skG_{2}+revG_{2}+\tau G_{2},U)\\ &=e(A.skG_{2}+revG_{2})\cdot e(\tau G_{2},U)\\ &=e(A.skG_{2},U)\cdot e(revG_{2},U)\cdot e(\tau G_{2},U)\\ &=e(G_{2},A.skU)\cdot e(G_{2},revU)\cdot e(G_{2},\tau U)\\ &=e(G_{2},A.skU)\cdot e(G_{2},revU+\tau U)\\ &=e(G_{2},A.skU+revU+\tau U)\\ &=e(G_{2},s)\\ \end{split}$
where each equality holds for the bilinearity of the Miller loop.
3.5 Revocation
To make sure no revocation has been emitted for each credential, we update the revocation list from the issuer. A revocation list can be publicly distributed since revocation keys do not provide any information on the identity of holders, unless they are presented in a verification.
Given an element $c_{p}=\{id,p,r,s\}$ of the credential proof table and the Issuer public key $A.pk$, if the claim id is present in the revocation list we can take the corresponding revocation private key $rev$ and compute the factor $\tau$ as
$\tau=\mathcal{H}(e(A.pk,r)^{rev})$ (3)
The equation (3) gives the same result as (1) since using the bilinearity of the pairing [12] we have:
$\begin{split}\tau&=\mathcal{H}(e(A.pk,r)^{rev})\\ &=\mathcal{H}(e(A.pk,G_{1}\cdot\varepsilon)^{rev})\\ &=\mathcal{H}(e(A.pk,G_{1})^{rev\cdot\varepsilon})\\ &=\mathcal{H}(e(A.pk,G_{1}\cdot rev)^{\varepsilon})\\ &=\mathcal{H}(e(A.pk,r_{1})^{\varepsilon})\end{split}$
Then we can verify if a credential claim has been revoked checking first if the revocation public key match:
$G_{2}\cdot rev=p-G_{2}\cdot\tau$
and then verify the following signature with the issuer’s public key (pk):
$sig=s-sign(\tau,id)-sign(rev,id)$ (4)
This signature should be verified with the issuer’s pk since:
[…]
then exploiting the homomorphic property of BLS signature we got that $sig$ in (4) is signed using the issuer’s private key.
3.6 Threshold Revocation
In order to split responsibilities over interactive revocation we introduce an optional threshold over the revocation key, plus we split the functionality of credential issuance from that of revocation issuance, now operated by different peers.
This is implemented via an interactive process facilitated by a third party trustee who should be:
  • never entitled to publish revocations
  • connected to credential issuers to complete any credential signature
  • never informed about the identity of credential holders
  • regularly connected to revocation issuers to distribute shares
Such a trustee will be in contact with credential issuers for the signature of credentials: it will create the $rev$ revocation key while concealing it from them. The trustee then proceeds creating the $\sigma_{rev}=rev\cdot U$ revocation signature and the $r_{1}=rev\cdot G_{1}$ projection as detailed in the issuance phase (3.2). Finally the revocation trustee will distribute $\sigma_{rev}$ and $r_{1}$ to issuers for further processing into the final signed credential.
The trustee will then proceed asynchronously to split the revocation key into shares using a public verifiable secret sharing (PVSS [13]) implementation and distribute these shares to all revocation issuers. In order to issue a revocation, a configurable quorum of peers among the revocation issuers will need to reconstruct the secret revocation key and publish it.
4 Benchmarks
We implemented the flows for credential issuance, presentation, verification and revocation for lab tests using Zenroom  [^1] Zenroom home: https://zenroom.org, a secure isolated execution environment implementing advanced cryptography transformations. The reference implementation for this paper is published on a public repository  [^2] SD-BLS github repository: https://github.com/dyne/sd-bls. All benchmarks were executed on a 6th gen. Intel PC running tests on a single i7 3.40GHz core and making no use of hardware acceleration.
Lab measurements on a growing number of claims show that issuance is less computation heavy and that the production of proofs and their verification need approximately the same resources, as shown in figure 1.
Figure 1: Speed comparison of issue, prove and verify
Based on our benchmarks, the resulting data objects sizes are:
  • Claim: 322 Bytes
  • Proof: 226 Bytes
  • Revocation: 64 Bytes
The computational cost grows for the verification phase with the presence of a cryptographic revocations list. We assume that revocations are grouped by claim type or credential structure, for instance all revocations of a certain claim (e.g.: "is above 18" or "is a Danish citizen") for any holder are published in the same list: this will not degrade the level of privacy and will group all the revocations in smaller lists.
Lab measurements of the time taken by a single proof verification process to operate on a growing number of revocations demonstrate that there is a linear growth, as shown in figure 2.
Figure 2: Speed of verification of a claim over multiple revocations
At last we measure the binary hamming distance between unlinkable proofs of the same credential signed by the same issuer. We performed this test without using any special random generator. This measurement should serve as a quick evaluation of the "randomness" of all blinded contents constituting a proof: the signature, the issuer’s public key and the revocation’s public key, from this benchmark the claim’s identifier is excluded. The results are shown in figure 3 and include an overlay of frequency values taken from the comparison of two random values. This benchmarks show that the binary hamming distance between different proofs is the same as the one between two random values of the same length generated by the same PRNG.
Figure 3: Frequency of hamming distance between unlinkable proofs
5 Conclusion
5.1 Security considerations
The revocations database is privacy and corruption sensitive (by our previous definition of issuer corruption) and it should be securely stored by each Issuer. This is mitigated by the optional adoption of threshold revocation. When using threshold revocation there is still the need for a revocation trustee who may be a single point of failure if allowed to publish revocation keys dishonestly collected during the process of credential issuance.
BLS signatures and the proof system obtained with credentials are considered secure by assuming the existence of random oracles [14], together with the decisional Diffie-Hellman Problem (DDH) [15], the external Diffie-Hellman Problem (XDH), and with the Lysyanskaya-Rivest-Sahai-Wol Problem (LRSW) [16], which are connected to the Discrete Logarithm. In fact, under these assumptions, we have that our protocol satisfies unforgeability, blindness, and unlinkability.
The SD-BLS implementation we are presenting in this paper is demonstrated using the BLS12-381 curve [11] also adopted by ETH2.0. Debating the choice of BLS12-381 is beyond the scope of this paper, but is worth mentioning that we can easily switch using the BLS461 curve based on a 461 bit prime, hence upgrading our implementation to 128 bit security [17] against attacks looking for discrete logs on elliptic curves [18].
The future growth of quantum-computing technologies may be able to overcome the Discrete Logarithmic assumptions by qualitatively different computational means and SD-BLS may then be vulnerable to quantum-computing attacks. However this is speculative reasoning on what we can expect from the future.
5.2 Future development
In this section we describe possibilities for expanding the algorithm to cover further applications, which appear promising while requiring further investigation.
Key-value claims
The focus of the initial version of the SD-BLS algorithm is on claims consisting of fixed and predetermined values (e.g.: "is above 18" or "has M.Sc. in applied cryptography"), whereby a credential could contain multiple claims. This initial version doesn’t include the possibility to process claims containing a predetermined part (the key) and a variable one (the value). We’re expecting to be able support key-value based claims in the next version of the paper.
Compatibility with EUDI-ARF
EUDI-ARF dictates that the holder’s secret key generation and signatures must occur inside a trusted platform module (TPM). For mobile devices, this limits the secret keys and signatures to those offered, via proprietary APIs by the mobile OS, namely RSA (multiple flavours) and ECDSA on the secp256r1 curve. Currently the TPMs APIs supported by Android and iOS do not support BLS 12381 key generation or signature. Client-side signatures in EUDI-ARF are mostly used in the authentication process, specifically in the proof of possession required by the OpenID4VCI[19] issuance flow, but not in the verification. Therefore, we can investigate the possibility to use SD-BLS to implement a partially retro-compatible superset of EUDI-ARF, by maintaining the current issuance and verification protocols and using an extended SD-JWT format. Also, because of the unlinkable feature of credentials and revocations in SD-BLS, we can explore useful integrations with the European Blockchain Services Infrastructure (EBSI [20]).
Digital Product Passport
Efforts in standardization of Digital Product Passport (DPP) are ongoing in both the EU (Cirpass, BatteryPass, Trace4EU) and US (DSCSA). An obstacle to adoption of DPP technologies is the reluctancy of manufacturers to share information about their supply-chain, knowing that the information would become publicly available and immutable due to blockchain storage. While requiring further analysis and investigation, a further development of the SD-BLS scheme could allow creating DPPs built on the selective disclosure principles, which may facilitate the adoption of the technology in the industry by preserving the privacy of natural persons present in the DPP as REA agents[21], while authenticating their contribution.
DAO Technologies
The SD-BLS math is fully compatible with ETH2.0 and can be computed inside an Ethereum VM. A verifier implemented in solidity can be a building block for more advanced Distributed Autonomous Organizations (DAO [22]) that want to authenticate peers using on a set of universally recognizable credentials instead of a key based proof of possession.

### v2

@@
 Abstract
-It is of critical importance to design digital identity systems that ensure the privacy of citizens as well protect them from issuer corruption.
-Unfortunately, what Europe’s and USA’s public sectors are currently developing does not offer such basic protections.
-Aiming at solving this issue, we propose a method for untraceable selective disclosure and privacy preserving revocation of digital credentials, utilizing the unique homomorphic characteristics of second order Elliptic Curves and Boneh-Lynn-Shacham (BLS) signatures operated on them.
-Our approach ensures that users can selectively reveal only the necessary credentials, while protecting their privacy across multiple presentations.
-We also aim to protect users from issuer corruption, by making it possible to apply a threshold for revocation to require collective agreement among multiple revocation issuers.
+It is of critical importance to design digital identity systems that ensure the privacy of citizens as well as protecting them from issuer corruption.
+We aim to solve this issue and propose a method for selective disclosure and privacy preserving revocation of digital credentials, using the unique homomorphic characteristics of second order Elliptic Curves and Boneh-Lynn-Shacham (BLS) signatures.
+Our approach ensures that users can selectively reveal credentials signed by a certain issuer, which can be interactively revoked by a quorum of other agreeing issuers without revealing the identity of users.
+Our goal is to protect users from issuer corruption by requiring collective agreement among multiple revocation issuers.
 1 Introduction
@@
 Credentials are signed by issuer authorities and encapsulated within various forms of digital proofs to be held in digital wallets, empowering individuals to reveal only chosen details to designated recipients, to limit data exposure and permit a user-controlled release of information.
-Such systems are known as selective disclosure and this article aims at improving their cryptography to adhere to basic privacy-by-design standards.
+Such systems are known as selective disclosure and they enhance users privacy by allowing data minimization when proving credentials.
 1.1 State of the art
 Selective disclosures are being used by nation states across the world in their next generation identity systems, for instance EIDAS2.0 in Europe where the European Digital Identity Wallet Architecture and Reference Framework[1] mandates the use of SD-JWT[2] and mDOC [3].
-Unfortunately both SD-JWT and mDOC adopt for their cryptography just Hash-Based Message Authentication Codes (HMAC) to generate proofs, and they can be easily traced (non-unlinkability): recent critical feedback on the EUDI ARF[4] details several problems.
-In North America the efforts concentrate on the adoption of the the BBS+ algorithm[5] leveraging its Zero Knowledge Proof properties and applied to W3C Verifiable Credentials[6] to obtain an higher degree of privacy.
+SD-JWT adopts for its cryptography Hash-Based Message Authentication Codes (HMAC) to generate proofs: such presentations can be traced and recent critical feedback on the EUDI ARF[4] details this problem.
+In North America the efforts concentrate on the adoption of the BBS+ algorithm[5] leveraging its Zero Knowledge Proof properties and applied to W3C Verifiable Credentials[6] to obtain an higher degree of privacy by making every disclosure unlinkable.
 1.2 Threats considered
@@
 Every verifiable presentation includes one or more HMAC(s), formatted in SD-JWT: the HMACs are identical each time a verifiable presentation is produced from a certain credential.
-This makes possible for colluding relying parties, or to malicious actors, to trace a holder’s identity by collecting, exchanging and confronting verifiable presentations.
-In order to preserve the privacy of a credential holder the disclosures presented by the wallet should not be traceable across different presentations (unlinkability).
+This makes possible for colluding relying parties, or to malicious actors, to trace a holder’s identity by collecting, exchanging and confronting verifiable presentations (linkability).
 This threat appears to be well mitigated by BBS+ through its Zero Knowledge Proof implementation.
 Privacy breach of revocation lists
+We believe that anonymous revocations are a condicio sine qua non to guarantee a sufficient level of privacy in digital identities and credentials.
 There is no privacy-preserving revocation system designed, either in EUDI-ARF, or in W3C-VC and BBS+.
-The approach followed for the revocation is that of the status (revocation) lists [7] for W3C-VC and EUDI-ARF is investigating a similar method.
-The use of a certificate status lists (CRL) presents issues related primarily to privacy [8], because sensitive information about holders leaks from the list.
-This is partially mitigated by expiration dates, in cases where credentials can be short-lived (typically less important credentials), but not applicable with digital identification documents such as ID, driving license, passport and social security numbers, which typically have longer or no expiration time.
 In case the choice of strategy for revocation is left open to developers, the risk for major privacy breaches may occur, for example with the adoption of public status lists [7].
-Based on these arguments, we believe that anonymous revocations are a sine qua none condition to guarantee a sufficient level of privacy in digital identities and credentials.
-Our approach is that of designing a privacy-preserving revocation mechanism to remove the leak of holder’s information.
+The hypotetical use of a certificate status lists (CRL) presents issues related primarily to privacy [8], because sensitive information about holders leaks from the list.
+This problem is partially mitigated by expiration dates, in cases where credentials can be short-lived (typically less important credentials), but not applicable with digital identification documents such as ID, driving license, passport and social security numbers, which typically have longer or no expiration time.
+Future plans for the national standards we are observing include the adoption of “Bitstring” status lists [9] which may grant a degree of privacy.
+Our approach is that of designing a privacy-preserving revocation mechanism to remove the leak of holder’s information and allow to delegate the governance of revocations to multiple revocation issuers which may be different from the credential issuer.
 Revocation issuer corruption
@@
 Similarly, a security breach of an issuer service, would result in similar threats.
-We mitigate this risk by introducing the possibility for threshold issuance of revocation keys.
+We mitigate this risk by introducing the possibility for threshold issuance of revocation keys and by separating the responsibility of revocation issuance and credential issuance.
+Feature Comparison
+UP UR TR
+SD-BLS no yes yes
+SD-JWT no wip no
+BBS+ yes wip no
+Table 1: feature comparison
+We briefly round up on feature differences between the named selective disclosure cryptographic schemes, as shown in table 1, mainly distinguishing between three fundamental features:
+  • UP: Unlinkable Presentation
+  • UR: Unlinkable Revocation
+  • TR: Threshold Revocation
+Where wip is mentioned meaning work in progress on adoption of bitstring status lists for unlinkable revocations.
 2 Overview
 2.1 Key contributions
-The cryptographic scheme described in the paper, name SD-BLS for brevity, takes inspiration by modern zero knowledge proofs schemes such as [5]BBS signatures or Coconut[10].
-The paper aims to employ some of the properties of the said schemes, overcome some of the shortcomings of SD-JWT and mDOC formats, mainly the non-unlinkability of credentials, proposing a novel cryptographic approach to similar data structures.
+The cryptographic scheme described in this paper, named SD-BLS for brevity, implements some of the properties of the SD-JWT scheme (which can be easily encoded in mDOC format) and proposes a novel cryptographic approach to similar data structures.
 Furthermore SD-BLS proposes novel anonymous cryptographic revocation flow for verifiable credentials, that aims to solve the privacy issues posed by status and revocation lists.
-Unlinkabe selective disclosure
+Selective disclosure
 Similarly to the SD-JWT and mDOC formats, SD-BLS produces an array of claims: the elements of the array are individually signed by the issuer.
 In SD-BLS the signature(s) replace the HMAC.
-Using the homomorphic properties of this curve SD-BLS enables the holder to produce a different presentation each time they want to prove a claim.
-SD-BLS can be used to prove credentials in public (on blockchain) when they are boolean attributes.
+Using the homomorphic properties of this curve SD-BLS enables the holder to selectively disclose only certain issuer signed claims and produce a presentation that minimizes private information given to verifiers.
 Anonymous cryptographic revocation
-SD-BLS proposes a novel approach to credential revocation: the data published by the revocation issuer will produce cryptographic material that contains no information about and is never linkable to the holder.
+SD-BLS proposes a novel approach to credential revocation: the data published by the revocation issuer will produce cryptographic material that contains no information about the credential holders.
 The cryptographic revocation material allows anyone to verify if an SD-BLS proof produced by a credential holder has been revoked by one or multiple issuers.
-The unlinkability, and thus anonymity of the cryptographic revocation, allows the revocation issuer to publish revocations off or on blockchain.
+The unlinkability, and thus anonymity of the cryptographic revocation, allows the revocation issuer to share revocations in public and allows anyone to verify if credentials have been revoked.
 2.2 Applications
-In this section we present some applications and use cases for digital identity and credentials, that could benefit from using the unlinkability and anonymous revocation of the the SD-BLS scheme.
+In this section we present some applications and use cases for digital identity and credentials, that could benefit from using the SD-BLS scheme.
 Digital identity
@@
 Similarly, the US government is experimenting with W3C-VC and mDOC for cross-states interoperable driving licenses.
-SD-BLS data format is similar to SD-JWT and mDOC, offering selective disclosure and adding unlinkability and anonymous revocation.
+SD-BLS data format is similar to SD-JWT and mDOC, offering selective disclosure and anonymous revocation.
 Academic credentials
@@
 We are unaware of standardization efforts for interoperable credentials in the fields of "Know Your Customer" (KYC) and Anti Money-Laundering (AML) certifications.
-We are aware of solution providers experimenting with W3C-VC for AML applications.
+We are aware of solution providers experimenting with W3C-VC for AML applications and believe SD-BLS can greatly improve the governance of credential revocation, which is a critical component for this use case.
 Generic light credentials
@@
 Verifiable credentials on Blockchain
-SD-BLS can be used on blockchain to produce publicly verified proofs when the credentials are made of clear-text messages stating a boolean property or ACL clearance.
+SD-BLS can be used with blockchain-based smart-contracts to activate certain functions:
   • A verifiable presentation can be used by a smart-contract to verify if one or more holder’s claim match the requirement needed to process a transaction.
-The verifiable presentation is anonymous, therefore its storage in the transaction log doesn’t pose privacy issues and is GDPR compatible.
-  • Issuers can publish their cryptographic revocation lists on chain, allowing smart-contracts to verify the status of a credential, still granting unlinkability.
+  • Issuers can publish their cryptographic revocation lists on chain, allowing smart-contracts to verify the status of a credential.
   • The issuer’s public keys can also be published on-chain, although this does not represent a novelty.
-It’s worth mentioning that, while the verifiable presentation is itself unlinkable to the holder’s blockchain address, the unlinkability of the holder to their blockchain address at verification phase, needs to be addressed within the blockchain protocol used: this issue is out of scope in this paper.
 3 Implementation
@@
 Since $\sigma_{i}\in G_{1}\forall i$, the algorithm has an homomorphic property. We exploit this property to add a revocation signature into the signed credential.
-We create a new secret key $rev$ with public key $r_{2}$, then compute the point $r_{1}\in\mathbb{G}_{1}$ and signature of the claim $\sigma_{rev}$ to add to the issuer signature:
-$\begin{split}r_{2}&=rev\cdot G_{2}\\ r_{1}&=rev\cdot G_{1}\\ \sigma_{rev}&=rev\cdot U\\ \sigma&=sk\cdot U+\sigma_{rev}\\ \end{split}$
+We create a new secret key $rev$ with public key $r$ and signature of the claim $\sigma_{rev}$ to add to the issuer signature:
+$\begin{split}r&=rev\cdot G_{2}\\ \sigma_{rev}&=rev\cdot U\\ \sigma&=sk\cdot U+\sigma_{rev}\\ \end{split}$
 The set of all the signed claims will be:
-$\mathcal{C}=\big{\{}\{r_{1i},r_{2i},\sigma_{i}\}:i\in claims\big{\}}$
+$\mathcal{C}=\big{\{}\{id,r,\sigma\}:id\in claims\big{\}}$
 At the end of this phase the holder is sent the signed claims to be stored in a private wallet, while the issuer stores a list of revocations, associated with an identifier of the holder and the specific credential, into a private database that can be used later to issue revocations.
 3.3 Presentation
-As we have done in the issuance phase (3.2), we exploit the homomorphic property of BLS signatures once again, but this time to render the presented signature unlinkable (impossible to trace across multiple presentations) by adding a "blinding factor" signature to the one provided by the issuer.
-Let $\mathcal{H}$ denote a cryptographic hash function (e.g.
-SHA-256) and let $\varepsilon$ be an integer chosen uniformly at random that serves as the blinding factor.
-Using the same technique of the Tripartite Diffie-Hellman Key Exchange [12] we construct a secret that is binded to the blinding factor, the Issuer public key and the "revocation point" $r_{1}=rev\cdot G_{1}$ computed in the issuance phase:
-$\tau=\mathcal{H}(e(A.pk,r_{1})^{\varepsilon})$ (1)
-When the holder needs to disclose a particular claim $m$ we construct the credential proof as follow: let $\sigma$ be the signature of the claim provided by the Issuer, then apply the blinding factor to the signature using the secret $\tau$
-$s=\sigma+sign(\tau,m)$ (2)
-and the corresponding public key computed as
-$p=r_{2}+G_{2}\cdot\tau$
-where $r_{2}$ is the revocation key linked to the claim. We also attach the public key $r=G_{1}\cdot\varepsilon$ that correspond to the secret key $\varepsilon$.
+When the holder needs to disclose a particular claim $id$ it is sufficient to send the signed claim set $\{id,r,\sigma\}$.
 3.4 Verification
-Credential verification is made by checking that the issuer has signed the blinded proof being presented. In order to do that the credential issuer’s public key must be added to the proof.
+Credential verification is made by checking the presented issuer’s signature. In order to do that the credential issuer’s public key must be added to the revocation public key.
 As explained in section 3.3 we can consider the credential proof as a collection of tables of the following form:
-$c_{p}=\{id,p,r,s\}$
-where $s$ and $p$ are respectively the signature and the public key of the credential proof of the string $id$ after verifying the status of the revocation, we can check the validity of the presented claim computing the key
-$pk=A.pk+p$
-and verify the bls signature.
+$c_{id}=\{id,r,\sigma\}$
+where $\sigma$ and $r$ are respectively the signature of the string $id$ and the revocation public key. We can check the validity of the presented claim computing the key
+$pk=A.pk+r$
+and verify the bls signature $\sigma$.
 3.4.1 Proof of correctness
-The signature $s$ is given by
-$\begin{split}s&=\sigma+sign(\tau,id)\quad\text{where}\\ \sigma&=sign(A.sk,id)+sign(rev,id)\\ \text{so}\quad s&=A.skU+revU+\tau U\end{split}$
+The signature $\sigma$ is given by
+$\begin{split}\sigma&=sign(A.sk,id)+sign(rev,id)\\ &=A.skU+revU\end{split}$
 where $U$ is the mapping of the string $id$ in the group $G_{1}$.
 Recalling the verification formula, it holds that:
-$\begin{split}e(pk,U)&=e(A.skG_{2}+revG_{2}+\tau G_{2},U)\\ &=e(A.skG_{2}+revG_{2})\cdot e(\tau G_{2},U)\\ &=e(A.skG_{2},U)\cdot e(revG_{2},U)\cdot e(\tau G_{2},U)\\ &=e(G_{2},A.skU)\cdot e(G_{2},revU)\cdot e(G_{2},\tau U)\\ &=e(G_{2},A.skU)\cdot e(G_{2},revU+\tau U)\\ &=e(G_{2},A.skU+revU+\tau U)\\ &=e(G_{2},s)\\ \end{split}$
+$\begin{split}e(pk,U)&=e(A.skG_{2}+revG_{2},U)\\ &=e(A.skG_{2},U)\cdot e(revG_{2},U)\\ &=e(G_{2},A.skU)\cdot e(G_{2},revU)\\ &=e(G_{2},A.skU+revU)\\ &=e(G_{2},\sigma)\\ \end{split}$
 where each equality holds for the bilinearity of the Miller loop.
+This concludes the first verification phase. If the given presentation is valid, then the verifier should proceed to check the revocation status of the credential as explained below.
 3.5 Revocation
-To make sure no revocation has been emitted for each credential, we update the revocation list from the issuer.
-A revocation list can be publicly distributed since revocation keys do not provide any information on the identity of holders, unless they are presented in a verification.
-Given an element $c_{p}=\{id,p,r,s\}$ of the credential proof table and the Issuer public key $A.pk$, if the claim id is present in the revocation list we can take the corresponding revocation private key $rev$ and compute the factor $\tau$ as
-$\tau=\mathcal{H}(e(A.pk,r)^{rev})$ (3)
-The equation (3) gives the same result as (1) since using the bilinearity of the pairing [12] we have:
-$\begin{split}\tau&=\mathcal{H}(e(A.pk,r)^{rev})\\ &=\mathcal{H}(e(A.pk,G_{1}\cdot\varepsilon)^{rev})\\ &=\mathcal{H}(e(A.pk,G_{1})^{rev\cdot\varepsilon})\\ &=\mathcal{H}(e(A.pk,G_{1}\cdot rev)^{\varepsilon})\\ &=\mathcal{H}(e(A.pk,r_{1})^{\varepsilon})\end{split}$
-Then we can verify if a credential claim has been revoked checking first if the revocation public key match:
-$G_{2}\cdot rev=p-G_{2}\cdot\tau$
-and then verify the following signature with the issuer’s public key (pk):
-$sig=s-sign(\tau,id)-sign(rev,id)$ (4)
-This signature should be verified with the issuer’s pk since:
-[…]
-then exploiting the homomorphic property of BLS signature we got that $sig$ in (4) is signed using the issuer’s private key.
+To control if a revocation has been emitted for any credential being verified, we update the revocation list from the issuer.
+A revocation list can be publicly distributed since revocation keys do not provide any information on the identity of holders.
+Given an element $c_{id}=\{id,r,\sigma\}$ of the credential presentation and the Issuer public key $A.pk$, if the claim id is present in the revocation list, we can take the corresponding revocation private key $rev$.
+We can verify the validity of the revocation by checking if the revocation public key presented by the holder match:
+$r=G_{2}\cdot rev$
+This step should follow the signature verification and is sufficient to guarantee the revocation status.
+If a dishonest holder provides a wrong $r$, then the verification of the signature will fail, and the given credential should be considered invalid.
+If it happens that the signature verification is successful and that the revocation key does not match, one can conclude that the credential is not revoked.
 3.6 Threshold Revocation
-In order to split responsibilities over interactive revocation we introduce an optional threshold over the revocation key, plus we split the functionality of credential issuance from that of revocation issuance, now operated by different peers.
-This is implemented via an interactive process facilitated by a third party trustee who should be:
+In order to split responsibilities over interactive revocation we introduce a threshold over the revocation key, plus we split the functionality of credential issuance from that of revocation issuance, now operated by different peers.
+This is implemented via an interactive process facilitated by a third party trustee, a revocation dealer, who should be:
   • never entitled to publish revocations
@@
   • regularly connected to revocation issuers to distribute shares
-Such a trustee will be in contact with credential issuers for the signature of credentials: it will create the $rev$ revocation key while concealing it from them.
-The trustee then proceeds creating the $\sigma_{rev}=rev\cdot U$ revocation signature and the $r_{1}=rev\cdot G_{1}$ projection as detailed in the issuance phase (3.2).
-Finally the revocation trustee will distribute $\sigma_{rev}$ and $r_{1}$ to issuers for further processing into the final signed credential.
-The trustee will then proceed asynchronously to split the revocation key into shares using a public verifiable secret sharing (PVSS [13]) implementation and distribute these shares to all revocation issuers.
+Such a revocation dealer will be in contact with credential issuers for the signature of credentials: it will create the $rev$ revocation key while concealing it from them.
+The trustee then proceeds creating the $\sigma_{rev}=rev\cdot U$ revocation signature and the $r=rev\cdot G_{2}$ public key as detailed in the issuance phase (3.2).
+Finally the revocation trustee will distribute $\sigma_{rev}$ and $r$ to issuers for further processing into the final signed credential.
+The revocation trustee (dealer) will then proceed asynchronously to split the revocation key into shares using a public verifiable secret sharing (PVSS [11]) implementation and distribute these shares to all revocation issuers.
 In order to issue a revocation, a configurable quorum of peers among the revocation issuers will need to reconstruct the secret revocation key and publish it.
+This process separates responsibilities between the credential issuer and the revocation issuers, delegating to the revocation issuers the possibility to revoke a credential interactively through a collective process.
+The collection of shares can be done asynchronously and is provable.
+The dealer should publish proofs of knowledge of each revocation share, proving their creation and authenticity.
+Revocation issuers can also use the dealer proofs to verify the validity of the shares received without revealing them.
+Such revocation proofs will also be useful when revocation issuers will reconstruct a revocation, since they can refer to them to prove their shares are authentic without revealing their content.
 4 Benchmarks
@@
 Intel PC running tests on a single i7 3.40GHz core and making no use of hardware acceleration.
-Lab measurements on a growing number of claims show that issuance is less computation heavy and that the production of proofs and their verification need approximately the same resources, as shown in figure 1.
-Figure 1: Speed comparison of issue, prove and verify
-Based on our benchmarks, the resulting data objects sizes are:
-  • Claim: 322 Bytes
-  • Proof: 226 Bytes
-  • Revocation: 64 Bytes
+Lab measurements on a growing number of claims show that issuance is less computation heavy than verification, as shown in figure 1.
+Figure 1: Speed comparison of issue and verify
+Based on our benchmarks, the resulting data objects sizes, excluding the signed material which can be either of variable length or fixed length hashes, are:
+  • Proof: 145 Bytes
+  • Revocation: 32 Bytes
 The computational cost grows for the verification phase with the presence of a cryptographic revocations list.
-We assume that revocations are grouped by claim type or credential structure, for instance all revocations of a certain claim (e.g.: "is above 18" or "is a Danish citizen") for any holder are published in the same list: this will not degrade the level of privacy and will group all the revocations in smaller lists.
+We assume that revocations are grouped by claim type or credential structure, for instance all revocations of a certain common claim (e.g.: "is above 18" or "is a Danish citizen") for any holder are published in the same list: this will not degrade the level of privacy and will group some revocations in smaller lists.
 Lab measurements of the time taken by a single proof verification process to operate on a growing number of revocations demonstrate that there is a linear growth, as shown in figure 2.
 Figure 2: Speed of verification of a claim over multiple revocations
-At last we measure the binary hamming distance between unlinkable proofs of the same credential signed by the same issuer.
-We performed this test without using any special random generator.
-This measurement should serve as a quick evaluation of the "randomness" of all blinded contents constituting a proof: the signature, the issuer’s public key and the revocation’s public key, from this benchmark the claim’s identifier is excluded.
-The results are shown in figure 3 and include an overlay of frequency values taken from the comparison of two random values.
-This benchmarks show that the binary hamming distance between different proofs is the same as the one between two random values of the same length generated by the same PRNG.
-Figure 3: Frequency of hamming distance between unlinkable proofs
+The threshold operated for revocation issuers consists of a verifiable secret sharing implementation supporting a configurable total and quorum of peers.
+Our implementation shows very good performance on reconstruction, which is the most speed-sensitive operation in scenarios where responsive revocation process is required.
+The process of reconstruction can be easily scaled for asynchronous consensus on a micro-service swarm architecture and verified on blockchain.
+Figure 3: Speed of creation and reconstruction of shares among multiple peers
 5 Conclusion
@@
 The revocations database is privacy and corruption sensitive (by our previous definition of issuer corruption) and it should be securely stored by each Issuer.
-This is mitigated by the optional adoption of threshold revocation.
+This is mitigated by the adoption of threshold revocation.
 When using threshold revocation there is still the need for a revocation trustee who may be a single point of failure if allowed to publish revocation keys dishonestly collected during the process of credential issuance.
 BLS signatures and the proof system obtained with credentials are considered secure by assuming the existence of random oracles [12], together with the decisional Diffie-Hellman Problem (DDH) [13], the external Diffie-Hellman Problem (XDH), and with the Lysyanskaya-Rivest-Sahai-Wol Problem (LRSW) [14], which are connected to the Discrete Logarithm.
-In fact, under these assumptions, we have that our protocol satisfies unforgeability, blindness, and unlinkability.
 The SD-BLS implementation we are presenting in this paper is demonstrated using the BLS12-381 curve [10] also adopted by ETH2.0.
@@
 However this is speculative reasoning on what we can expect from the future.
+The SD-BLS implementation we present in this paper doesn’t integrate any protection against replay-attacks, nor implements timestamps and expiration.
+We believe security features protecting from replay-attacks can be easily added and adapted from the context of deployment.
+I absence of session management providing such protections, any real-world use of SD-BLS should contemplate adding a signed timestamp to presentations and expiration checks on verifications.
 5.2 Future development
 In this section we describe possibilities for expanding the algorithm to cover further applications, which appear promising while requiring further investigation.
-Key-value claims
-The focus of the initial version of the SD-BLS algorithm is on claims consisting of fixed and predetermined values (e.g.: "is above 18" or "has M.Sc.
-in applied cryptography"), whereby a credential could contain multiple claims.
-This initial version doesn’t include the possibility to process claims containing a predetermined part (the key) and a variable one (the value).
-We’re expecting to be able support key-value based claims in the next version of the paper.
 Compatibility with EUDI-ARF
@@
 Therefore, we can investigate the possibility to use SD-BLS to implement a partially retro-compatible superset of EUDI-ARF, by maintaining the current issuance and verification protocols and using an extended SD-JWT format.
-Also, because of the unlinkable feature of credentials and revocations in SD-BLS, we can explore useful integrations with the European Blockchain Services Infrastructure (EBSI [20]).
+Also we can explore useful integrations with the European Blockchain Services Infrastructure (EBSI [18]).
+Signroom and DIDroom
+In SD-BLS both credential issuers and revocation issuers are in charge of various interactive administrative operations, while the dealing of revocation shares can be easily automated.
+We plan to integrate SD-BLS in the free and open source software “Signroom”, an application we developed in the context of the NGI ASSURE grant, and “DIDroom” the dashboard connected to our did:dyne W3C DID domain.
 Digital Product Passport
@@
 The SD-BLS math is fully compatible with ETH2.0 and can be computed inside an Ethereum VM.
-A verifier implemented in solidity can be a building block for more advanced Distributed Autonomous Organizations (DAO [22]) that want to authenticate peers using on a set of universally recognizable credentials instead of a key based proof of possession.
+A verifier implemented in solidity can be a building block for more advanced Distributed Autonomous Organizations (DAO [20]) that want to authenticate peers using the selective disclosure of verifiable credentials instead of a key based proof of possession.
+6 Acknowledgements
+This work has been funded by the EU in the framework of the NGI TRUSTCHAIN project, grant No 101093274.
+We thank Puria Nafisi Azizi and Matteo Cristino for their infatiguable work on Zenroom, Luca Di Domenico for his help on the PVSS implementation, Giuseppe De Marco for sharing findings from his pioneering journey, Simone Onofri for his insights on threat models and Gabriele Bellini for his helpful security review.
+We extend our thanks the anonymous reviewers for their valuable advice and the De Cifris association for facilitating connections over a large network of unique professionals.

### v4

@@
 Abstract
-It is of critical importance to design digital identity systems that ensure the privacy of citizens as well as protecting them from issuer corruption.
-We aim to solve this issue and propose a method for selective disclosure and privacy preserving revocation of digital credentials, using the unique homomorphic characteristics of second order Elliptic Curves and Boneh-Lynn-Shacham (BLS) signatures.
-Our approach ensures that users can selectively reveal credentials signed by a certain issuer, which can be interactively revoked by a quorum of other agreeing issuers without revealing the identity of users.
-Our goal is to protect users from issuer corruption by requiring collective agreement among multiple revocation issuers.
+Ensuring privacy and protection from issuer corruption in digital identity systems is crucial.
+We propose a method for selective disclosure and privacy-preserving revocation of digital credentials using second-order Elliptic Curves and Boneh-Lynn-Shacham (BLS) signatures.
+We make holders able to present proofs of possession of selected credentials without disclosing them, and we protect their presentations from replay attacks.
+Revocations may be distributed among multiple revocation issuers using publicly verifiable secret sharing (PVSS) and activated only by configurable consensus, ensuring robust protection against issuer corruption.
+Our system’s unique design enables fast revocation checks allowing the use of large revocation lists.
 1 Introduction
@@
 Future plans for the national standards we are observing include the adoption of “Bitstring” status lists [9] which may grant a degree of privacy.
-Our approach is that of designing a privacy-preserving revocation mechanism to remove the leak of holder’s information and allow to delegate the governance of revocations to multiple revocation issuers which may be different from the credential issuer.
+In SD-BLS we design a privacy-preserving revocation mechanism to remove the leak of holder’s information and delegate the governance of revocations to a quorum of multiple revocation issuers which may be different from the credential issuer.
 Revocation issuer corruption
@@
 We mitigate this risk by introducing the possibility for threshold issuance of revocation keys and by separating the responsibility of revocation issuance and credential issuance.
-Feature Comparison
-UP UR TR
-SD-BLS no yes yes
-SD-JWT no wip no
-BBS+ yes wip no
+2 Overview
+2.1 Feature Comparison
+UP UR TR URG
+SD-BLS no yes yes yes
+SD-JWT no wip no no
+BBS+ yes wip no yes
 Table 1: feature comparison
-We briefly round up on feature differences between the named selective disclosure cryptographic schemes, as shown in table 1, mainly distinguishing between three fundamental features:
+We briefly round up on feature differences between the named selective disclosure cryptographic schemes, as shown in table 1, mainly distinguishing between four fundamental features:
   • UP: Unlinkable Presentation
@@
   • TR: Threshold Revocation
-Where wip is mentioned meaning work in progress on adoption of bitstring status lists for unlinkable revocations.
-2 Overview
+  • URG: Unregroupability
+Where wip is mentioned, it means work in progress on adoption of bitstring status lists for unlinkable revocations.
+The “unregroupability” feature refers to the fact that different presentations of different claims cannot be linked to each other (regrouped) as presented by the same holder, even when they are signed by the same issuer.
 2.2 Key contributions
-The cryptographic scheme described in this paper, named SD-BLS for brevity, implements some of the properties of the SD-JWT scheme (which can be easily encoded in mDOC format) and proposes a novel cryptographic approach to similar data structures.
-Furthermore SD-BLS proposes novel anonymous cryptographic revocation flow for verifiable credentials, that aims to solve the privacy issues posed by status and revocation lists.
+The cryptographic scheme described in this paper, named SD-BLS for brevity, implements all the properties of the SD-JWT scheme and proposes a novel cryptographic approach to similar data structures.
+Furthermore SD-BLS proposes novel anonymous cryptographic revocation flow for verifiable credentials, that aims to solve governance issues posed by status and revocation lists.
 Selective disclosure
 Similarly to the SD-JWT and mDOC formats, SD-BLS produces an array of claims: the elements of the array are individually signed by the issuer.
-In SD-BLS the signature(s) replace the HMAC.
-Using the homomorphic properties of this curve SD-BLS enables the holder to selectively disclose only certain issuer signed claims and produce a presentation that minimizes private information given to verifiers.
+In SD-BLS the signature(s) replace the HMAC and still enable the holder to selectively disclose only certain signed credentials, and produce a proof of possession that minimizes private information given to verifiers.
 Anonymous cryptographic revocation
 SD-BLS proposes a novel approach to credential revocation: the data published by the revocation issuer will produce cryptographic material that contains no information about the credential holders.
-The cryptographic revocation material allows anyone to verify if an SD-BLS proof produced by a credential holder has been revoked by one or multiple issuers.
+The cryptographic revocation material allows anyone to verify if an SD-BLS proof produced by a credential holder has been revoked.
 The unlinkability, and thus anonymity of the cryptographic revocation, allows the revocation issuer to share revocations in public and allows anyone to verify if credentials have been revoked.
+Multi-stakeholder governance of revocations
+SD-BLS allows to introduce a new trusted party to the issuance phase: the revocation dealer.
+The dealer doesn’t need to know the content of the credential or the identity of an holder: its role is that of producing the revocation signature and distributing its secret key to a configurable range of revocation issuers.
+Later on a configurable quorum of issuers may reconstruct the revocation secret key to revoke a credential.
+This way the decision on a revocation is not delegated to a single peer, but to a multi-stakeholder governance that protects the holder from issuer corruption.
+Fast revocation checks
+SD-BLS shows good benchmark results on revocation checks: the main operation a verifier needs to do is an additional $\mathbb{G}_{2}$ multiplication.
+The cost of this operation grows linearly with the number of published revocations.
 2.3 Applications
@@
 Digital identity
-The focus of the EUDI-ARF specifications is identity documents: it defines mechanisms and data structures to issue a Personal Identification (PID) as well as digital driving licenses.
+The focus of the EUDI-ARF specifications is identity documents: it defines mechanisms and data structures to issue a Personal Identification (PID) as, for instance, with digital driving licenses.
 Similarly, the US government is experimenting with W3C-VC and mDOC for cross-states interoperable driving licenses.
-SD-BLS data format is similar to SD-JWT and mDOC, offering selective disclosure and anonymous revocation.
+The SD-BLS data format is similar to SD-JWT and mDOC, offering selective disclosure and anonymous revocation.
 Academic credentials
@@
 SD-BLS can be used with blockchain-based smart-contracts to activate certain functions:
-  • A verifiable presentation can be used by a smart-contract to verify if one or more holder’s claim match the requirement needed to process a transaction.
+  • A proof of possession can be published and be peer-verified on-chain in its private form without disclosing the credential contents.
+  • A smart-contract may verify if one or more holder’s credentials match the requirement needed to process a transaction.
   • Issuers can publish their cryptographic revocation lists on chain, allowing smart-contracts to verify the status of a credential.
@@
   • $E_{T}$ represents instead the group of points of the twisted curve of BLS12-381, with embedding degree $k=12$. The order of this group is the same of that of $E$;
-We also require defining the notion of a cryptographic pairing.
-That is defined as a function $e:\mathbb{G}_{2}\times\mathbb{G}_{1}\to\mathbb{G}_{T}$, where $\mathbb{G}_{1},\mathbb{G}_{2}$ and $\mathbb{G}_{T}$ are all groups of same order $n$, such that satisfies the following properties:
+We also require the notion of a cryptographic pairing. [11]
+For the purpose of our protocol we will consider the Miller pairing $e:E_{T}\times E\to\mathbb{G}_{T}$, where $\mathbb{G}_{T}\subset\mathbb{F}_{p^{12}}$ is the subgroup containing the $n$-th roots of unity, and $n$ is the order of the groups $E$ and $E_{T}$.
+For completeness we also recall the main properties of the map:
   • i.
-    Bilinearity, i.e. given $P_{1},Q_{1}\in\mathbb{G}_{1}$ and $P_{2},Q_{2}\in\mathbb{G}_{2}$, we have
+    Bilinearity, i.e. given $P_{1},Q_{1}\in E_{T}$ and $P_{2},Q_{2}\in E$, we have
     $e(P_{2},P_{1}+Q_{1})=e(P_{2},P_{1})\cdot e(P_{2},Q_{1})$
@@
   • ii.
-    Non-degeneracy, meaning that for all $g_{1}\in\mathbb{G}_{1},g_{2}\in\mathbb{G}_{2}$, $e(g_{2},g_{1})\neq 1_{\mathbb{G}_{T}}$, the identity element of the group $\mathbb{G}_{T}$;
+    Non-degeneracy, meaning that for all $g_{1}\in E_{T},g_{2}\in E$, $e(g_{2},g_{1})\neq 1_{\mathbb{G}_{T}}$, the identity element of the group $\mathbb{G}_{T}$;
   • iii.
@@
   • iv.
-    $\mathbb{G}_{1}\neq\mathbb{G}_{2}$, and moreover, that there exist no efficient homomorphism between $\mathbb{G}_{1}$ and $\mathbb{G}_{2}$.
-For the purpose of our protocol we will have $\mathbb{G}_{1}=E$ and $\mathbb{G}_{2}=E_{T}$, and $\mathbb{G}_{T}\subset\mathbb{F}_{p^{12}}$ is the subgroup containing the $n$-th roots of unity, where $n$ is the order of the groups $E$ and $E_{T}$.
-Instead $e:E_{T}\times E\to\mathbb{G}_{T}$ is the Miller pairing, which in our work is encoded as the method miller(ECP2 P, ECP Q).
+    $E_{T}\neq E$, and moreover, that there exist no efficient homomorphism between $E_{T}$ and $E$.
 3.2 Issuance
-As for other well known algorithms BLS signing will work following three main steps:
+As for other well known algorithms BLS signing works following three main steps:
   • Key Generation phase.
@@
 The corresponding public key $pk$ is the element $sk\cdot G_{2}\in E_{T}$;
-  • Signing phase. The credential $m$ is first hashed into the point $U\in E$, which in our scheme is done by the method hashtopoint; the related signature is then given by $\sigma=sk\cdot U$;
-  • Verification phase. For an other user that wants to verify the authenticity and the integrity of the message $m$, it needs to
+  • Signing phase.
+    The credential $m$ is first hashed into the point $U\in E$; the related signature is then given by $\sigma=sk\cdot U$;
+  • Verification phase.
+    For an other user that wants to verify the authenticity and the integrity of the message $m$, it needs to
     • 1.
@@
     If verification passes it means that $\sigma$ is a valid signature for $m$.
-  • Proof of correctness By using the definitions of the elements involved and exploiting the property of the pairing $e$ we have
-    $\begin{split}e(pk,U)&=e(sk\cdot G_{2},U)\\ &=e(G_{2},U)^{sk}\\ &=e(G_{2},sk\cdot U)\\ &=e(G_{2},\sigma)\end{split}$
 BLS signatures also support aggregation: it is possible to aggregate a collection of multiple signatures $\sigma_{i}$ (each one related to a different message $m_{i}$) into a singular new object $\sigma$, that can be validated using the respective public keys $pk_{i}$ in a suitable way.
 Since $\sigma_{i}\in G_{1}\forall i$, the algorithm has an homomorphic property. We exploit this property to add a revocation signature into the signed credential.
-We create a new secret key $rev$ with public key $r$ and signature of the claim $\sigma_{rev}$ to add to the issuer signature:
-$\begin{split}r&=rev\cdot G_{2}\\ \sigma_{rev}&=rev\cdot U\\ \sigma&=sk\cdot U+\sigma_{rev}\\ \end{split}$
+The issuer create a credential as follow: given a claim $m$, it generate a new secret revocation key $rev$ with public key $r$. Let $\mathcal{H}$ be a cryptographic hash function, we compute:
+$H=\mathcal{H}(m:r)$
+then the issuer proceeds to sign with both its private key, and the revocation key generated above:
+$\begin{split}r&=rev\cdot G_{2}\\ \sigma_{rev}&=sign(rev,H:r)\\ \sigma&=sign(sk,H:r)+\sigma_{rev}\\ \end{split}$
 The set of all the signed claims will be:
-$\mathcal{C}=\big{\{}\{id,r,\sigma\}:id\in claims\big{\}}$
-At the end of this phase the holder is sent the signed claims to be stored in a private wallet, while the issuer stores a list of revocations, associated with an identifier of the holder and the specific credential, into a private database that can be used later to issue revocations.
+$\mathcal{C}=\big{\{}\{H,r,\sigma,m\}:m\in claims\big{\}}$
+At the end of this phase the holder is sent the signed claims to be stored in a private wallet, while the issuer stores a list of revocations as tuples formed by $\{H,rev\}$ into a private database that can be used later to issue revocations.
 3.3 Presentation
-When the holder needs to disclose a particular claim $id$ it is sufficient to send the signed claim set $\{id,r,\sigma\}$.
+Any presentation of a SD-BLS credential can simply omit the message $m$ to separate disclosure from verification as required by the specific context of a proof of possession and to satisfy privacy preserving design patterns for data minimization.
+Basic Proof
+A credential holder can choose any set of signed claims to present, and selectively disclose them into what we name “Basic Proof” presentation.
+The holder can present a basic proof of possession $\{H,r,\sigma\}$ for each claim requested, or the complete set $\{H,r,\sigma,m\}$ for claims whose content must be disclosed.
+However, Both forms of presentation of a basic proof are vulnerable to replay attacks: any verifier receiving such presentations can reuse them to impersonate the holder in any other session.
+One Time Proof
+To prevent replay attacks we exploit once again the homomorphic property of BLS signatures to add a sessions signature into the credential signature, then we add separate fields containing a signed timestamp and the public key of the session signature.
+We name this presentation “One Time Proof”.
+For each presentation the holder generates a key pair $(sk_{t},pk_{t})$ and a string $t$ containing session information, i.e.
+an expiration date or a pointer to the intended recipient.
+Then we compute:
+$\begin{split}\sigma^{\prime}&=\sigma+sign(sk_{t},H:r)\\ \sigma_{t}&=sign(sk_{t},H:r:\sigma^{\prime}:t:pk_{t})\end{split}$
+And obtain a presentation composed as follows:
+$\{H,r,\sigma^{\prime},t,pk_{t},\sigma_{t}\}$
+This schema is secure because it is impossible to reconstruct the holder information $\{H,r,\sigma\}$ necessary for a reply attack; indeed, given $\sigma^{\prime}$, an attacker that wants to retrieve $\sigma$ should be able to reconstruct the signature $\sigma_{t}$, but this is not possible without the knowledge of the secret key $sk_{t}$.
+Furthermore the three attributes $t,\sigma_{t},pk_{t}$ cannot be modified or removed because $pk_{t}$ is necessary for the verification of signature $\sigma^{\prime}$.
+If an attacker tries to create a new valid presentation with freshly generated $sk_{t}^{\prime},pk_{t}^{\prime},t^{\prime}$ and $\sigma_{t}^{\prime}$, then the signature $\sigma^{\prime\prime}=\sigma^{\prime}+sign(sk_{t}^{\prime},H:r)$ can be verified using the public key $pk_{t}+pk_{t}^{\prime}$.
+However this public key does not verify the signature $\sigma_{t}^{\prime}$, and $\sigma_{t}^{\prime}$ can not be updated using the homomorphic property since it is the signature over the old $\sigma^{\prime}$.
+Thus any tampering of the presentation will lead to an invalid credential.
 3.4 Verification
-Credential verification is made by checking the presented issuer’s signature. In order to do that the credential issuer’s public key must be added to the revocation public key.
-As explained in section 3.3 we can consider the credential proof as a collection of tables of the following form:
-$c_{id}=\{id,r,\sigma\}$
-where $\sigma$ and $r$ are respectively the signature of the string $id$ and the revocation public key. We can check the validity of the presented claim computing the key
+Credential verification is made by checking the presented issuer’s signature and revocation status of each claim.
+In order to verify the signature the credential issuer’s public key must be added to the revocation public key.
+Basic proof
+We consider the basic credential proof as a collection of tables of the following form:
+$c=\{H,r,\sigma\}$
+where $\sigma$ and $r$ are respectively the signature of the string $H$ and the revocation public key, optionally the claim value $m$ can be included.
+We can check the validity of the presented claim computing the key:
 $pk=A.pk+r$
 and verify the bls signature $\sigma$.
-3.4.1 Proof of correctness
-The signature $\sigma$ is given by
-$\begin{split}\sigma&=sign(A.sk,id)+sign(rev,id)\\ &=A.skU+revU\end{split}$
-where $U$ is the mapping of the string $id$ in the group $G_{1}$.
+As proof of correctness consider that the signature $\sigma$ is given by:
+$\begin{split}\sigma&=sign(A.sk,H:r)+sign(rev,H:r)\\ &=A.skU+revU\end{split}$
+where $U$ is the mapping of the string $H$ in the group $G_{1}$.
 Recalling the verification formula, it holds that:
@@
 where each equality holds for the bilinearity of the Miller loop.
-This concludes the first verification phase. If the given presentation is valid, then the verifier should proceed to check the revocation status of the credential as explained below.
+When the presentation contains the value $m$ it should also be checked that:
+$H=\mathcal{H}(m:r).$
+One Time Proof
+If the verifier receives the credential locked to the current session (with replay attack protection) as $\{H,r,\sigma^{\prime},t,pk_{t},\sigma_{t}\}$ then it proceeds to
+  • verifiy the information in $t$ (e.g. timestamp)
+  • verify $\sigma_{t}$ with public key $pk_{t}$
+  • verify $\sigma^{\prime}$ with the public key $A.pk+r+pk_{t}$
+In case the verifier is presented with a one time disclosure that includes $m$ then it needs also to check its hashed value $H=\mathcal{H}(m:r)$.
+This concludes the first verification phase. If the given presentation is valid, then the verifier should proceed to check the revocation status of the credential.
 3.5 Revocation
@@
 A revocation list can be publicly distributed since revocation keys do not provide any information on the identity of holders.
-Given an element $c_{id}=\{id,r,\sigma\}$ of the credential presentation and the Issuer public key $A.pk$, if the claim id is present in the revocation list, we can take the corresponding revocation private key $rev$.
-We can verify the validity of the revocation by checking if the revocation public key presented by the holder match:
+Given an element $c=\{H,r,\sigma\}$ of the credential presentation, and a series of $rev$ elements found in any revocation list, we can verify a revocation by checking if the revocation public key presented by the holder matches:
 $r=G_{2}\cdot rev$
@@
 If it happens that the signature verification is successful and that the revocation key does not match, one can conclude that the credential is not revoked.
+Note that, in the case the verifier received a credential in the “One Time Proof” form, the check for the revocation status does not change.
 3.6 Threshold Revocation
@@
 Such a revocation dealer will be in contact with credential issuers for the signature of credentials: it will create the $rev$ revocation key while concealing it from them.
-The trustee then proceeds creating the $\sigma_{rev}=rev\cdot U$ revocation signature and the $r=rev\cdot G_{2}$ public key as detailed in the issuance phase (3.2).
-Finally the revocation trustee will distribute $\sigma_{rev}$ and $r$ to issuers for further processing into the final signed credential.
-The revocation trustee (dealer) will then proceed asynchronously to split the revocation key into shares using a public verifiable secret sharing (PVSS [11]) implementation and distribute these shares to all revocation issuers.
+The dealer then proceeds creating the $\sigma_{rev}=sign(rev,H:r)$ revocation signature and the $r=rev\cdot G_{2}$ public key, which will be communicated to the credential issuer to be aggregated into the signed credential (see section (3.2)).
+The dealer will then proceed to split the secret revocation key into shares using a public verifiable secret sharing (PVSS [12]) implementation and distribute these shares to all revocation issuers.
 In order to issue a revocation, a configurable quorum of peers among the revocation issuers will need to reconstruct the secret revocation key and publish it.
@@
 Lab measurements on a growing number of claims show that issuance is less computation heavy than verification, as shown in figure 1.
-Figure 1: Speed comparison of issue and verify
-Based on our benchmarks, the resulting data objects sizes, excluding the signed material which can be either of variable length or fixed length hashes, are:
-  • Proof: 145 Bytes
+Figure 1: Speed comparison of issue and verify in SD-BLS
+As a handy reference, we provide a benchmark of the same operations performed using the BBS+ in Zenroom on the same machine: results show that SD-BLS is one order of magnitude slower than BBS+, whose verification runs aproximately at the same speed as other operations.
+Figure 2: Speed comparison of issue and verify in BBS+
+Based on our benchmarks, the resulting data objects sizes in SD-BLS are:
+  • Signed claim: 177 Bytes
+  • Proof: 322 Bytes
   • Revocation: 32 Bytes
-The computational cost grows for the verification phase with the presence of a cryptographic revocations list.
-We assume that revocations are grouped by claim type or credential structure, for instance all revocations of a certain common claim (e.g.: "is above 18" or "is a Danish citizen") for any holder are published in the same list: this will not degrade the level of privacy and will group some revocations in smaller lists.
-Lab measurements of the time taken by a single proof verification process to operate on a growing number of revocations demonstrate that there is a linear growth, as shown in figure 2.
+The computational cost of verifications grows linearly in presence of a cryptographic revocations list.
+We assume revocations are published as lists of secret revocation keys.
+Lab measurements of the time taken by a single proof verification process to operate on a growing number of revocations is shown in figure 3.
 Figure 3: Speed of verification of a claim over multiple revocations
@@
 This is mitigated by the adoption of threshold revocation.
-When using threshold revocation there is still the need for a revocation trustee who may be a single point of failure if allowed to publish revocation keys dishonestly collected during the process of credential issuance.
+When using threshold revocation there is still the need for a revocation dealer who has no access to private informations, but may be dishonest and publish revocation keys collected during the process of credential issuance.
 BLS signatures and the proof system obtained with credentials are considered secure by assuming the existence of random oracles [13], together with the decisional Diffie-Hellman Problem (DDH) [14], the external Diffie-Hellman Problem (XDH), and with the Lysyanskaya-Rivest-Sahai-Wol Problem (LRSW) [15], which are connected to the Discrete Logarithm.
+The future growth of quantum-computing technologies may be able to overcome the Discrete Logarithmic assumptions by qualitatively different computational means and SD-BLS may then be vulnerable to quantum-computing attacks.
+However this is speculative reasoning on what we can expect from the future.
 The SD-BLS implementation we are presenting in this paper is demonstrated using the BLS12-381 curve [10] also adopted by ETH2.0.
 Debating the choice of BLS12-381 is beyond the scope of this paper, but is worth mentioning that we can easily switch using the BLS461 curve based on a 461 bit prime, hence upgrading our implementation to 128 bit security [16] against attacks looking for discrete logs on elliptic curves [17].
-The future growth of quantum-computing technologies may be able to overcome the Discrete Logarithmic assumptions by qualitatively different computational means and SD-BLS may then be vulnerable to quantum-computing attacks.
-However this is speculative reasoning on what we can expect from the future.
-The SD-BLS implementation we present in this paper doesn’t integrate any protection against replay-attacks, nor implements timestamps and expiration.
-We believe security features protecting from replay-attacks can be easily added and adapted from the context of deployment.
-I absence of session management providing such protections, any real-world use of SD-BLS should contemplate adding a signed timestamp to presentations and expiration checks on verifications.
+The $H$ component may be protected against brute-forcing attacks using hash collisions to forge a valid credential for a different message.
+A protection against this attack can be the adoption of a key-derivation function like Argon2 [18] on hash creation, which will add computational costs to issuance and verification.
+Anyone with knowledge of $H,r$ can try to guess $m$ by appending known strings, which becomes trivial especially in case of boolean credential strings (i.e.
+“above18=true”).
+We recommend the credential issuer appends a nonce like $m:nonce$ before hashing and signing the message, which will be known by the holder and always disclosed when disclosing $m$.
 5.2 Future development
@@
 For mobile devices, this limits the secret keys and signatures to those offered, via proprietary APIs by the mobile OS, namely RSA (multiple flavours) and ECDSA on the secp256r1 curve.
-Currently the TPMs APIs supported by Android and iOS do not support BLS 12381 key generation or signature.
+Currently the TPMs APIs supported by Android and iOS do not support BLS 12-381 key generation or signature.
 Client-side signatures in EUDI-ARF are mostly used in the authentication process, specifically in the proof of possession required by the OpenID4VCI[19] issuance flow, but not in the verification.
@@
 The SD-BLS math is fully compatible with ETH2.0 and can be computed inside an Ethereum VM.
-A verifier implemented in solidity can be a building block for more advanced Distributed Autonomous Organizations (DAO [20]) that want to authenticate peers using the selective disclosure of verifiable credentials instead of a key based proof of possession.
+A verifier implemented in solidity can be a building block for more advanced Distributed Autonomous Organizations (DAO [22]) that want to authenticate peers using the selective disclosure of verifiable credentials instead of a single key based proof of possession.
 6 Acknowledgements
 This work has been funded by the EU in the framework of the NGI TRUSTCHAIN project, grant No 101093274.
-We thank Puria Nafisi Azizi and Matteo Cristino for their infatiguable work on Zenroom, Luca Di Domenico for his help on the PVSS implementation, Giuseppe De Marco for sharing findings from his pioneering journey, Simone Onofri for his insights on threat models and Gabriele Bellini for his helpful security review.
+We thank Puria Nafisi Azizi and Matteo Cristino for their infatiguable work on Zenroom, Luca Di Domenico for his help on the PVSS implementation, Giuseppe De Marco for sharing findings from his pioneering journey and Simone Onofri for his insights on threat models.
 We extend our thanks the anonymous reviewers for their valuable advice and the De Cifris association for facilitating connections over a large network of unique professionals.